Existence and Regularity for Mixturs of Micromagnetic Materials

E. Acerbi and Giuseppe Mingione Dipartimento di Matematica Viale Delle Scienze 43100 Parma, Italy acerbi@unipr.it mingione@unipr.it

Irene Fonseca Department of Mathematical Sciences Carnegie Mellon University Pittsburgh, PA 15213 fonseca@andrew.cmu.edu

Abstract. A new model for the energy of a mixture of micromagnetic materials is introduced within the context of functions with special bounded variation (SBV). Existence and regularity for the solution of an optimal design problem in micromagnetics are obtained.